OpenStudy (anonymous):

Solve for x: 1/(x^2+3x)-3/(x^2+2x-3)=1/(x^2-x)

OpenStudy (anonymous):

\[\frac{1}{x^2+3x}-\frac{3}{x^2+2x-3}=\frac{1}{x^2-x}\]factor the denominators \[\frac{1}{x(x+3)}-\frac{3}{(x+3)(x-1)}=\frac{1}{x(x-1)}\] multiply both sides by the least common multiple \(x(x-1)(x+3)\)to clear the fractions

OpenStudy (anonymous):

you should get \[x-1-3x=x+3\] which is relatively easy to solve

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yw